OpenStudy (anonymous):

2^-4/3^-2 Express the number in a/b form where a and b are integers

OpenStudy (anonymous):

start with \[\frac{3^2}{2^4}\] and compute

OpenStudy (anonymous):

idk how to when its negative

OpenStudy (anonymous):

flip them

OpenStudy (anonymous):

that makes the exponents positive, so you can compute

OpenStudy (anonymous):

\[\frac{2^{-4}}{3^{-2}}=\frac{3^2}{2^4}\] and both \(3^2\) and \(2^4\) are easy to compute